• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Ramesh/Xiaoyan

    Log entry time 18:12:35 on September 10, 2009

    Entry number 289004

    keyword=access activity (pvdis work)

    Checked both left and right HRS pvdis setup Dx10 signals. They look like
    other BCM signals. In left HRS, all BCM signals (including UNSER) are now
    perfectly digital (no attenuation any more) since they are re-routed from
    755 modules instead of a LeCroy level translator.

    We also found that in the right HRS pvdis low threshold discriminator,
    almost no channels gave clock output. Found that a too low threshold was
    the culprit. When set the threshold as -25 mV, we saw expected output of
    an input. Now the thresholds of low to high threshold discriminator are
    like this:
    low: -25 mV
    medium: -30 mV
    hig: -105 mV

    Checked left HRS TDC signals on the scope. All pulser signals are exactly
    750 ns earlier than the common stop (also provided by the pulser trigger)
    signal. Looks normal.